OpenStudy (anonymous):

A carpet layer just finished laying 228 sq ft of carpet in a square room. Now she must put a metal strip along one side of the carpet. Estimate the length of metal strip that she will need. Round your answer to the nearest whole number.

OpenStudy (anonymous):

From the given area, solve for the measurement of a side. Use formula for getting area of a square

OpenStudy (austinl):

\[A = l^2\] A = 228 \[228 = l^2\] \[l = \sqrt{228}\] That is one side.